john needed money for college. he borrowed $7500 at 6% simple interest per year. if he paid $180 interest, what was the duration of the loan.
Answer by fractalier(6550)

The formula for finding simple interest is
I = prt
Here we are looking for t, in years, so
t = I / pr = 180 / (7500)(.06) = 0.4 years